OpenStudy (anonymous):

Find the equation of the line that is perpendicular to the line formed by the equation 3x + 2y = -14 and passes through the point (-48, 3).

OpenStudy (anonymous):

First make this in y=mx+b form

OpenStudy (anonymous):

Then find the slope (m) and make it negative for the perpendicular line then, use point slope form with that new slope and build your new line equation y-y1 = m(x-x1)

OpenStudy (anonymous):

m = -3/2, so -m = 3/2

OpenStudy (anonymous):

so y-(3) = 3/2 * (x-(-48)

OpenStudy (anonymous):

y-3 = 3/2 * (x + 48)

OpenStudy (anonymous):

distribute the 3/2, then add 3 to both sides

OpenStudy (anonymous):

y = - 2/3 x - 29 correct

OpenStudy (anonymous):

\[y = \frac{3}{2}x + (\frac{3}{2}\times 48) + 3\]

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

\[y = \frac{3}{2} + 72 + 3\] \[y = \frac{3}{2} + 75\]

OpenStudy (anonymous):

perpendicular to the original equation

OpenStudy (anonymous):

going through that point (-48, 3)

OpenStudy (anonymous):

x+75 you mean

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[y=\frac{3}{2}x + 75\]

OpenStudy (anonymous):

medal me

OpenStudy (anonymous):

haa I have more questions but I just need you to correct me ...how do I medal you by the way

OpenStudy (anonymous):

click best

OpenStudy (anonymous):

and start a new question for each question

OpenStudy (anonymous):

it seems like you are having trouble building your initial equation by rearranging the terms given you need to isolate y in this equation: 3x + 2y = -14 first you subtract 3x from both sides giving 2y = -3x -14 then you divide by 2, both sides giving y = (-3/2)x - 7

OpenStudy (anonymous):

the coefficient on the x is your slope, m. in this case, it is -3/2. perpendicular lines have the slope with the opposite sign. so, perpendicular to this is 3/2

OpenStudy (anonymous):

we took that slope value and plugged it into the point-slope equation: y-y1 = m(x-x1) giving y-y1 = (3/2)(x-x1) where x1 and y1 and the x and y values from your point

OpenStudy (anonymous):

ya get it?

OpenStudy (anonymous):

now click "best answer"